OpenStudy (anonymous):

solve the exponential equation using the method of relating the bases by first rewriting the equation in the form of e^u=e^v 1/e ^6x =√e ÷e ^6−x

OpenStudy (anonymous):

again, -6x=1/2 +x-6 5.5=7x x=11/14

OpenStudy (anonymous):

thank you!!!